If the average life time of an excited state of H atom is of order 10-8 sec, estimate how many orbits an e- makes when it is in the state n=2 and before it suffers a transition to n=1 state


Grade:

1 Answers

vikas askiitian expert
13 years ago

speed of electron in any orbit = vn = vo/n                    (vo = 2.18*106m/s)

so , v2 = vo/2  = 1.09*106m/s

radius of  orbit = rn = aon2                      (ao = 52.9 * 10-12m)

so , r2 = 2.12*10-10m

 

v = X/T

x = @r      so                               ( r is radius of second orbit)

v = @r/T                                           ( @ is angular  displacement)

@ = vT/r

@ = 2pi(n)                                (here n is number of revolutions)

2pi(n) = vT/r        .......................1

substituting values of v,r,T we get

n = 8.18*106 revolutions

 

approve if u like this ans